OpenStudy (anonymous):

4i(1+i)+4 / 2i 2x^2 + 7z less than or greater to 4...put this one on a number line

OpenStudy (anonymous):

can someone plz help

OpenStudy (anonymous):

looks like two questions right? i am not sure i understand the second one

OpenStudy (anonymous):

well can you do number 1

OpenStudy (anonymous):

or the first one for that matter is it \[\frac{4i(1+i)+4}{2i}\]??

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

Thats it

OpenStudy (anonymous):

first multiply out in the numerator, get \[\frac{4i-4+4}{2i}\] or \[\frac{4i}{2i}\] then divide